x^2-16
x^2-6x+8
these are two different Questions
x^2-10x+25
x-5

Answer by venugopalramana(3286)

WHAT DO YOU WANT FACTORISATION?PLEAE WRITE QUESTION CORRECTLY
x^2-16 =(X)^2-(4)^2=(X+4)(X-4)...USING FORMULA A^2-B^2=(A+B)(A-B)
x^2-6x+8 =X^2-4X-2X+8=X(X-4)-2(X-4)=(X-4)(X-2)
these are two different Questions
x^2-10x+25=X^2-2*X*5+5^2=(X-5)^2...USING FORMULA A^2-2*A*B+B^2=(A-B)^2
x-5=X-5 ONLY ...IF YOU WANT DIVISION THEN
x^2-10x+25.....X^2-2*X*5+5^2...............(X-5)^2......X-5
--------- = -------------...............=--------- = ---- IS THE ANSWER
x-5.................X-5......................X-5..........1